OpenStudy (logan13):

What is the value of w? 5(9 − 3) = w • 9 – 5 • 3 w =

OpenStudy (logan13):

ryan got any ideas

OpenStudy (ryanwhite499):

Let's solve your equation step-by-step. 5(9−3)=w(9)−(5)(3) Step 1: Simplify both sides of the equation. 5(9−3)=w(9)−(5)(3) Simplify: (Show steps) 30=9w−15 Step 2: Flip the equation. 9w−15=30 Step 3: Add 15 to both sides. 9w−15+15=30+15 9w=45 Step 4: Divide both sides by 9. 9w 9 = 45 9 w=5 Answer: w=5

OpenStudy (logan13):

thanks ryan i got more if you want medals